OpenStudy (secret-ninja):

Medals. Which equation is NOT equivalent to the formula d = rt? r = t / d r = d / t t = d / r d = tr

OpenStudy (some.random.cool.kid):

oh ik this

OpenStudy (some.random.cool.kid):

i just did this one

OpenStudy (secret-ninja):

I think it MIGHT be B...

OpenStudy (anonymous):

it is A

OpenStudy (some.random.cool.kid):

it should be D or B

OpenStudy (some.random.cool.kid):

no not A

OpenStudy (some.random.cool.kid):

i believe its D

OpenStudy (some.random.cool.kid):

try D because i just did this one

OpenStudy (anonymous):

Valid forms of d=rt r=d/t t=d/r Nothing else works when singling out a variable.

OpenStudy (some.random.cool.kid):

its D

OpenStudy (some.random.cool.kid):

your welcome

OpenStudy (some.random.cool.kid):

bye

OpenStudy (anonymous):

D works, distance = rate x time.

OpenStudy (anonymous):

it is b because it is the formula just a different way to do this type of problem

OpenStudy (anonymous):

They just flipped the order of the variables.

OpenStudy (anonymous):

B works because rate is depending on the distance you travel over time.

OpenStudy (secret-ninja):

So it is D?

OpenStudy (anonymous):

It' is NOT D did is a valid equation.

OpenStudy (anonymous):

it is b

OpenStudy (secret-ninja):

Oh okay, so I was right? :D

OpenStudy (anonymous):

b

OpenStudy (anonymous):

yes you were good job

OpenStudy (anonymous):

I hope everyone understands she is looking for an INVALID equation.

OpenStudy (anonymous):

114 no she ain't you shut up

OpenStudy (anonymous):

"Which equation is NOT equivalent to the formula d = rt" emphasis on not equivalent.

OpenStudy (secret-ninja):

..... Erm.. its getting kinda tense here, I'll close the question... O.o

OpenStudy (anonymous):

If you really need help instead of getting confused by everyone else here message this, this is just insane. The question is being made more difficult than it should be because of all th enonesense.

OpenStudy (anonymous):

message me* not message this >.<

OpenStudy (anonymous):

looooooosssssseeeeeeeeeerrrrrrrrrrrrrrrrssssssssssss